Hello I need the formula to solve this problem:
Find the area and perimeter
of a square with a diagonal of 10 inches.

Answer by checkley77(12844)

A SQUARE WITH A DIAGONAL IS
X^2+X^2=10^2
2X^2=100
X^2=100/2
X62=50
X=SQRT50
X=7.07 IN. IS THE LENGTH OF THE SIDES OF THE SQUARE.
AREA=7.07^2=50 IN^2.
PERIMETER=4*7.0728.28 IN.